OpenStudy (gabby1214):

solve a-9=-17 the solution is a=?(Integer or simplified fraction)

OpenStudy (anonymous):

to solve this you will simply add -17+9=?

OpenStudy (anonymous):

a-9=-17 +9=-8 -8-9=-17 a=-8